OpenStudy (yinshy):

A bag contains 60 marbles of various colors. Shana draws 48 times, each time replacing the marble before drawing again. She gets a total of 8 green marbles. Which is best prediction of how many green marbles are in the bag? A. 4 B. 6 C. 8 D. 10

OpenStudy (yinshy):

b?

OpenStudy (anonymous):

What you think?

OpenStudy (yinshy):

B

OpenStudy (anonymous):

I'm not too sure about this one, but I believe it is C

OpenStudy (yinshy):

c? hmmm

OpenStudy (anthonyym):

Yes I think it's C too. @Yinshy, how did you get B?

OpenStudy (yinshy):

it sounded rigght

OpenStudy (anthonyym):

Any idea how to solve it?

OpenStudy (anthonyym):

Ok. So out of 48 times, she draws 8 marbles out of those 48 times. So you can get a ratio from that.

OpenStudy (anonymous):

\[8/48 = g/60\]g = 10 or D.

OpenStudy (anonymous):

90 % sure.